OpenStudy (anonymous):

Describe the distributive property? How is it different from other properties? Can you give an example?

OpenStudy (anonymous):

@texaschic101

OpenStudy (youngstudier):

The Distributive Property of Multiplication over addition states that instead of simplifying what is in parentheses first, you may choose to multiply each term by the factor outside the parentheses: a (b + c) = ab + ac

OpenStudy (youngstudier):

It's the same thing for subtraction : a( b - c) = ab - ac

OpenStudy (youngstudier):

Does that help at all @geny55 ?

OpenStudy (anonymous):

yes